OpenStudy (yanasidlinskiy):

A rectangular parking lot has a length that is 15 yards greater than the width. The area of the parking lot is 450 square yards. Find the length and the width. Use the formula, are = length * width.

OpenStudy (yanasidlinskiy):

15 yards greater than the width = x(x+15) Is that to what it translates to?

OpenStudy (johnweldon1993):

Let Width = W and Length = W + 15  If the area in total = 450 We have \[\large Length \times Width = 450\] \[\large (W+15) \times W = 450\] Solve for 'w' and then solve for 'l'